OpenStudy (anonymous):

problem: 3√54 / 3√-2 using radicals

OpenStudy (anonymous):

Same thing again, you need to make sure you write \[\sqrt[3]{54} / \sqrt[3]{-2}\] \[=\sqrt[3]{-27} = -3\]

OpenStudy (anonymous):

3√54 / 3√-2 = 3√[(54/(-2)] = 3√-27 = -3